OpenStudy (anonymous):

A 5-mile cab ride costs $7.20. A 9-mile cab ride costs $11.60. Find a linear equation that models a relationship between cost c and distance d. c = 1.44d + 4.40 d = 1.10c + 4.40 c = 1.29d + 1.70 c = 1.10d + 1.70

OpenStudy (anonymous):

@gigi.antonia

OpenStudy (anonymous):

D

OpenStudy (anonymous):

graph the set of data draw a trend line and write its equation {(1,7),(-2,1),(3,13),(-4,-3),(0,5)} A. y=0.9x+8.6 B. y=2/5x-2 C. y=5/2x+5 D. y=5

OpenStudy (anonymous):

number 5

OpenStudy (anonymous):

number 2

OpenStudy (anonymous):

@JohnnyDPug @jasonjohnson86 @justnick09 @kl0723 @maurice2232 @armine